OpenStudy (lilsis76):

using addition formulas for sine and cosine to simplify the expression: cos (2pi/9) cos (pi/18) + sin (2pi/9) sin(pi/18) I want to learn it the normal way first before I add my instructors instructions to place S=45 & T = 30. here are the formulas given: sin(s+t) = sin s cos t + cos s sine t sin(s-t) = sin s cos t - cos s sine t cos(s+t) = cos s cos t - sin s sin t cos(s-t) = cos s cos t + sin s sin t The they way it looks, it seems to use formula 4: cos(s-t) = cos s cos t + sin s sin t

OpenStudy (anonymous):

yes.... it is that 4th formula..

OpenStudy (lilsis76):

YAY! okay um... i know that know would i fill it in like this... well wait, first I have a question, how can i find out if its cos or sin that im looking for?

OpenStudy (lilsis76):

wait it would be the cos. right? ugh these formulas are confusing me

OpenStudy (anonymous):

yes... the cosine one... just remember, the cosine formula (whether it's + or -), is cos * cos +/- sin * sin

OpenStudy (lilsis76):

okay, ill have to put that on a 3by5 card

OpenStudy (lilsis76):

okay! i think i got this just a second. let me know if its right or i made a mistake please.

OpenStudy (anonymous):

so: \(\large cos(\frac{2\pi}{9}-\frac{\pi}{18})=cos(\frac{2\pi}{9})cos(\frac{\pi}{18})+sin(\frac{2\pi}{9})sin(\frac{\pi}{18}) \)

OpenStudy (lilsis76):

Cos(2pi/9 – pi/18) mult. 2 to the left side……. So cos(4pi/18-pi/18) I guess giving me…….cos 3pi/18? Or do I simplify to get the cos pi/6 ???

OpenStudy (anonymous):

now... \(\large \frac{2\pi}{9}-\frac{\pi}{18}=\frac{4\pi}{18}-\frac{\pi}{18}=\frac{3\pi}{18}=\frac{\pi}{6} \) that expression actually is cos(pi/6)

OpenStudy (anonymous):

are you familiar with the unit circle?

OpenStudy (lilsis76):

yes, i had to memorize that thingy

OpenStudy (anonymous):

wow... you seem like you don't need any help.... :)

OpenStudy (lilsis76):

so...my answer is going to be that unit?

OpenStudy (lilsis76):

ugh?! i do need h

OpenStudy (anonymous):

so what's cos(pi/6) = ???

OpenStudy (lilsis76):

haha sorry about that. i just need help of steps walking through it, she went so fast in class. i only got pieces

OpenStudy (lilsis76):

30deg

OpenStudy (anonymous):

yes... cos(pi/6) = cos(30 degrees) = ????

OpenStudy (lilsis76):

the points? is that what the ??? is for?

OpenStudy (anonymous):

yes.... what's the x-coordinate ??? here is the unit circle...page 3 http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (lilsis76):

\[\sqrt{3}/2 , 1/2\]

OpenStudy (anonymous):

i have to leave... sorry... the answer is cos(30 degrees) = \(\large \frac{\sqrt3}{2} \)

OpenStudy (lilsis76):

okay, thank u, bye!

OpenStudy (anonymous):

the x-coordinate refers to the cosine, the y-coordinate is sine. yw... :)